# Interpretable Concept Bottlenecks to Align Reinforcement Learning Agents

### v2

@@ 1 Introduction
Reinforcement learning (RL) agents, particularly deep RL agents are prone to suffer from a variety of issues that hinder them from learning optimal or generalisable policies. Prominent examples are reward sparsity (Andrychowicz et al., 2017) and difficult credit assignment (Raposo et al., 2021; Wu et al., 2023). A more recent, yet pressing issue is the goal misalignment problem. It occurs when an agent optimizes a different side-goal, that is aligned with the original target goal during training (Koch et al., 2021), but not at test time. Such misalignments can be difficult to identify (di Langosco et al., 2022). For instance in this work we discover for the first time that such a misalignment can occur in the one of the oldest and most iconic Atari games, Pong (cf. Fig. 1, top). In Pong, the agent’s target goal is to catch a ball with its own paddle and to return it passed the enemy’s one (Mnih et al., 2013). Due to the automated enemy, that constantly follows the ball, agents can learn to focus on the position of the enemy paddle’s for placing their own paddle, rather than the position of the ball itself. This phenomenon is more generally described by the umbrella term shortcut learning (Geirhos et al., 2020) in which a model learns decision rules for the provided training set, that fail to transfer to novel data sets. If left unchecked, it may lead to a lack of model generalization and dangerous and unintuitive failures when such models are deployed in real-world settings (Advani & Saxe, 2017; Zhang et al., 2021).
@@
Figure 1: Successive Concept Bottlenecks Agents (SCoBots) allow for easy inspection and revision. Top: Deep RL agents trained on Pong produce high playing scores with importance map explanations that suggest sensible underlying reasons of taking an action ( Refer to caption ). However, when the enemy is hidden, the deep RL agent fails to even catch the ball without clear reasons ( Refer to caption ). Bottom: SCoBots, on the other hand, allow for multi-level inspection of the reasoning behind the action selection , e.g., at a relational concept, but also action level. Moreover, they allow for users to easily intervene on them ( Refer to caption ) to prevent the agents from focusing on potentially misleading concepts. In this way SCoBots can mitigate RL specific caveats like goal misalignment.
@@
A recently popular approach for identifying such shortcut behaviour is via methods from the field of eXplainable AI (XAI), whose aim is to identify the reasons behind a model’s decisions (Lapuschkin et al., 2019; Guidotti et al., 2019; Schramowski et al., 2020; Ras et al., 2022; Roy et al., 2022; Saeed & Omlin, 2023). Many methods of XAI however suffer from issues such as a lack of faithfulness, i.e. the explanation does not faithfully present the model’s underlying decision process (Hooker et al., 2019; Chan et al., 2022; DeYoung et al., 2020), and coarse, low-level explanation semantics. For example, importance-map explanations indicate the importance of an input element for a decision without indicating why this element is important (Kambhampati et al., 2021; Stammer et al., 2021; Teso et al., 2023). A recent branch of research therefore focuses on developing models that provide inherent concept-based explanations. Prominent examples of this idea are concept bottlenecks models (CBMs) which have been shown to provide predictive performances on par with standard deep learning approaches for supervised image classification (Koh et al., 2020). More importantly, CBMs allow to identify and effectively revise incorrect model behaviour on a higher, concept level (Stammer et al., 2021).
Despite the success and advantageous properties of CBMs for tackling RL-specific issues discussed above, no research has however been performed for their integration into RL. A reason for this is that previous concept-based models depend only on concepts that are at a single object level, e.g., object color, shape, etc. However, objects in RL environments, even as simple as Atari games, interact with one another (Kaiser et al., 2019). In order to adequately solve such RL tasks via explicit concept representations, agents thus need to also extract necessary relational concepts.
@@
In this work we therefore introduce Successive Concept Bottleneck Agents (SCoBots, cf. Fig. 1, bottom) as an important extension of the original concept bottleneck approach to RL. SCoBots integrate successive concept bottlenecks into their decision processes where each bottleneck layer provides concept representations that integrate the concept representations of the previous bottleneck layer. Specifically, provided a set of predefined concept functions SCoBots automatically extract relational concept representations based on objects and their properties of the initial bottleneck layers. Finally, the set of relational and object concept representations are used for optimal action selection. SCoBots thus represent inherently explainable RL agents and in comparison to deep RL agents, they allow for inspecting and revising their learned decision policies at multiple levels of their reasoning processes: from the single object properties, through the relational concepts to the actual action decision.
@@
In our evaluations on the iconic Atari Learning Environments (Mnih et al., 2013) (ALE), we provide experimental evidence of the performance abilities of SCoBots that is on par with deep RL agents. More importantly, we showcase their ability to provide valuable explanations and the potential of mitigating a variety of RL specific issues, from reward sparsity to misalignment problems, via simple guidance from domain experts. In the course of our evaluations we identify previously unknown shortcut behaviour of deep RL agents, even on a simple game as Pong. By utilizing the interaction capabilities of SCoBots, this behaviour is easily removed. Ultimately, our work thereby illustrates the severity of goal misalignment issues in RL and the importance of being able to mitigate these and other RL specific issues via relational concept based models.
In summary our specific contributions are:
We introduce Successive Concept Bottleneck agents (SCoBots) that integrate relational concept representations into their decision processes.
@@
We show that SCoBots allow to inspect their internal decision processes.
@@
We show that the inherent inspectable nature of SCoBots can be utilized for identifying previously unknown misalignment problems of RL agents.
@@
We show that they allow for human interactions for mitigating various RL specific learning issues.
@@
We proceed as follows. We first formally introduce SCoBots and discuss their specific properties. We continue with experimental evaluations via several metrics and analysis. Before concluding, we touch upon related work.
@@
Figure 2: An overview of Successive Concept Bottlenecks Agents (SCoBots). SCoBots decompose the policy into consecutive interpretable concept bottlenecks (ICB). Objects and their properties are first extracted from the raw input, human-understandable functions are then employed to derive relational concepts and finally used to select an action. The understandable concepts enable interactivity. Each bottleneck allows expert users to, e.g., prune or utilize concepts to define additional reward signals.
@@ 2 Successive Concept Bottleneck Agents
Let us start with providing the background notations of RL problems, modelled as markov decision processes (MDPs), and also of concept bottleneck models (CBMs). In this work, we represent an RL problem through the framework of a Markov Decision Process, $\mathcal{M}=<\mathcal{S},\mathcal{A},P_{s,a},R_{s,a}>$, with $\mathcal{S}$ as the state space, $\mathcal{A}$ the set of available actions, $P(s,a)$ the transition probability, and $R(s,a)$ the immediate reward function, obtained from the environment. Classic deep RL policies, $\pi_{\theta}(s)=P(A=a|S=s)$ parametrized by $\theta$, are usually black-box models that process raw input states, e.g. a set of frames, to provide a distribution over the action space (Mnih et al., 2015; van Hasselt et al., 2016).
@@
The main body of research on CBMs focuses on image classification tasks, where extracted concepts represent attributes of objects. In contrast, RL agents must learn not just from static data but also through interaction with dynamic environments. RL tasks thus often require relational reasoning, as they involve understanding the relationships between instances that evolve through time and interact with another. This is crucial for learning effective policies in complex, dynamic environments. Note that, in the following descriptions, we use specific fonts to distinguish objects’ properties from their relations.
@@ 2.1 Building inspectable ScoBots
An underlying assumption of SCoBots is that their processing steps should be inspectable and understandable by a human user. This stands in contrast to e.g. unsupervised CBM approaches (Jabri et al., 2019; Zhou et al., 2019; Srinivas et al., 2020), where there is no guarantee for learning human-aligned concepts. SCoBots rather take a different approach by dividing the concept learning and RL-specific action selection into several inspectable steps that are grounded in human-understandable concept representations. These steps are described hereafter and depicted in Fig. 2.
@@
One of the major differences to previous work on CBMs is that SCoBots further extract and utilize relational concepts that are based on the previously extracted objects and their properties. Formally, SCoBots utilize a consecutive bottleneck model, $\mu_{\mathcal{F}}(\cdot)$, called the relation extractor. This model, $\mu$, is parameterized by a predetermined set of transparent relational functions, $\mathcal{F}$, used to extract a set of $d_{t}$ relational concepts. These relations are based on each individual object (and its properties) for unary relations or on combinations of objects for n-ary relations and are denoted as $\mu_{\mathcal{F}}(\Omega_{t})=\Gamma_{t}=\{g_{t}^{k}\}_{k=1}^{d_{t}}$. Without strong prior knowledge, $\mathcal{F}$ can initially correspond to universal object relations such as distance and speed. However, this set can easily be updated on the fly by a human user with e.g. additional, novel relational functions. Note that $\mathcal{F}$ can include the identity, to let the relation extractor pass through initial object concepts from $\Omega_{t}$ to the relational concepts $\Gamma_{t}$.
@@
Finally, SCoBots employ an action selector, $\rho_{\theta_{2}}$, parameterized by $\theta_{2}$}, on the relational concepts to select the best action, $a_{t}$, given the initial state, $s_{t}$ (i.e. the set of frames). Up to now, all extracted concepts in SCoBots, both $\Omega_{t}$ and $\Gamma_{t}$, represent human-understandable concept representations. To guarantee understandability also within the action selection step of SCoBots, we need to embed an interpretable action selector. While neural networks, a standard choice in RL literature, are performative and easy to train using differentiable optimization methods, they lack this required interpretability feature. However, (Çağlar Aytekin, 2022) have recently shown the equivalence between ReLU-based neural networks and decision trees, which in contrast represent inherently interpretable models. To trade-off these issues of flexibility and performance vs interpretability, SCoBots thus break down the action selection process by initially training a small ReLU-based neural network action selector, $\widetilde{\rho}_{\theta^{\prime}_{2}}$, via gradient-based RL optimization. After this, $\widetilde{\rho}_{\theta^{\prime}_{2}}$ is finally distilled into a decision tree $\rho_{\theta_{2}}$. Lastly, note that one can add a residual link from the initial object concepts, $\Omega_{t}$, to the relational concepts, $\Gamma_{t}$ such that the action selector can, if necessary, also make decisions based on basic object properties, e.g., the height of an object.
Overall, our approach preserves the MDP formulation used by classic deep approaches, but decomposes the classic deep policy (1) into a successive concept bottleneck one (2):
$s_{t}$ $\xrightarrow{\pi_{\theta}}a_{t}$ (1)
$s_{t}\xrightarrow{\omega_{\theta_{1}}}\Omega_{t}$ $\xrightarrow{\mu_{\mathcal{F}}}\Gamma_{t}\xrightarrow{\rho_{\theta_{2}}}a_{t}$ (2)
@@
where $\theta=(\theta_{1},\theta_{2},\mathcal{F}$) constitutes the set of policy parameters. For simplicity, we will discard the parameter notations in the rest of the manuscript. Further explanations of the input and output space of each module, as well as the properties and relational functions used in this work are provided in the appendix (cf. App.A.5 and App. A.6).
@@
Instead of jointly learning object detection, concept extraction, and policy search, SCoBots enable the optimization of each policy component independently. Separating the training procedure of different components (similar to that of (Koh et al., 2020)) particularly reduces the complexity of the overall optimization problem.
@@ 2.2 Guiding SCoBots
Concept pruning. The type and amount of concepts that are required for a successful policy may vary greatly across tasks. For instance, the objects colors are irrelevant in Pong but required to distinguish vulnerable ghosts from dangerous ones in MsPacman. At the same time overloading the action selector with too many irrelevant concepts, whether these are object properties ($\Omega_{t}$) or relational concepts ($\Gamma_{t}$), can lead to difficult optimization (e.g. the agent focusing on noise in unimportant states) as well as difficult inspection of the decision tree ($\rho$). Moreover, for a single RL task, the need for specific concepts might even change during training, e.g. in progressive environments (where agents are required to master early stages before being provided with additional, more complex tasks (Delfosse et al., 2024)).
SCoBot’s human-comprehensible concepts therefore allow domain experts to prune unnecessary concepts. In this way, expert users can guide the learning process towards relevant concepts. Formally, users can (i) select a subset of the object property concepts, $\overline{\Omega}$. Additionally, by (ii) selecting a subset of the relational functions, $\overline{\mathcal{F}}$, or (iii) specifying which objects specific functions should be applied on, experts can implicitly define the relational concepts subset, $\overline{\Gamma}$. Guided SCoBots thus formally refine Eq. 2, leading to:
@@
Furthermore, users can (iv) prune out redundant actions resulting in a new action space $\overline{\mathcal{A}}$.
Figure 3: Object-centric agents can master different Atari environments and interactive SCoBots allow for corrections. Human-normalized scores of deep and SCoBots (with and without guidance) agents trained using PPO on 99 ALE environments. SCoBots obtain similar or better scores than the deep agents, showing that object-centric agents can also solve RL tasks while making use of human-understandable concepts (left). Guiding SCoBots allow to correct misalignment in Pong (center) and to obtain the originally intended agents, depicted by a level complition score of 100% on the intended goal’s evaluation in Kangaroo (right).
@@
For example, experts have the ability to impose penalties, based on the distance between the agent and objects (cf. “expert reward” arrow in Fig. 2), with the intention of incentivizing the agent to maintain close proximity with the ball. As we show in our experimental evaluation, we can use this concept pruning and concept based reward shaping to easily address many RL specific caveats such as reward sparsity, ill-defined objectives, difficult credit assignment, and misalignment (cf. App. A.8 for details on each problem).
@@ 3 Experimental Evaluations
In our evaluations, we investigate several properties and potential benefits of the SCoBot agents. We specifically aim to answer the following research questions:
@@
(Q2) Does the inherent inspectable nature of SCoBots allow to detect potential issues in their decision processes?
@@
(Q3) Can concept-based guidance help mitigate RL caveats of SCoBot agents, such as policy misalignments?
Figure 4: SCoBots allow to follow their decision process, because of their object centricity and interpretable concepts. The decision processes of SCoBots (extracted from the decision trees) with the associated states from a guided SCoBot on Skiing (left), and from normal SCoBots on Pong (middle) and Kangaroo (right). For example, in this Skiing state, our SCoBot selects RIGHT, as the signed distance between Player and the (left) Flag1 (on the $x$ axis) is bigger than $15$. This agent selects the correct action for the right reasons.
Experimental setup: We evaluate SCoBots on $9$ Atari games (cf. Fig. 3 (left) for complete list) from the Atari Learning Environments (Bellemare et al., 2012) (by far the most used RL framework (cf. App. A.1)), as well as a variation of Pong where the enemy is not visible yet active in the game (i.e. still able to return the ball).
@@
For results based on achieved game scores we provide human normalized scores (following eq. 6) that are averaged over $3$ seeds for each agent configuration. We compare our SCoBot agents to deep agents with the classic convolutional network introduced by (Mnih et al., 2015), that process a stack of $4$ black and white frames and denote these as deep agents in the following. Note, that we evaluate all agents on the latest v5 version of the environments [^2] This leads to deep agents overall slightly worse than in (Schulman et al., 2017), however we obtain similar results when evaluating on the old v4 versions (cf. Tab. 1). to prevent overfitting. All agents are trained for $20\text{M}$ frames under the Proximal Policy Optimization algorithm (PPO, (Schulman et al., 2017)), specifically the stable-baseline3 implementation (Raffin et al., 2021) and its default hyperparameters (cf. Tab. 2 in App. A.5).
@@
We focus our SCoBot evaluations on the more interesting aspects of the agent’s reasoning process underlying the action selection, rather than the basic object identification step (which has been thoroughly investigated in previous works e.g. (Koh et al., 2020; Stammer et al., 2021; Wu et al., 2023)). We thus assume access to a pretrained object extractor and provide our agents with object-centric descriptions of these states ($\Omega_{t}$) based on the information from OCAtari (Delfosse et al., 2023a). Specifically, in our evaluations, the set of object properties consists of object class (e.g. enemy paddle, ball etc.), $(x^{t},y^{t})$ coordinates, coordinates at the previous position $(x^{t-1},y^{t-1})$, height and width, the most common RGB values (i.e. the most representative color of an object), and object orientation. The specific set of functions, $\mathcal{F}$, that are used to extract object relations is composed of: euclidean distance, directed distances (on $x$ and $y$ axes), speed, velocity, plan intersection, the center (of the shortest line between two objects), and color name (cf. App. A.5.1 for implementation details).
@@
To distill SCoBots’ learned policies, we use a rule-based model extraction method, similar to (Zarlenga et al., 2021). For the human-guided SCoBot evaluations (denoted as (guided) SCoBots in our evaluations), we illustrate human guidance and prune out the concepts that we consider unimportant to master the game (cf. Tab. 4). Furthermore, to mitigate RL specific problems, we also provide SCoBot agents with simulated human-expert feedback signals, the details of which we describe at the relevant sections below. More details about the experimental setup and the hyperparameters are provided in App. A.5.
@@
The main target of developing ScoBots is to obtain competitive, yet transparent agents that allow for human users to identify their underlying reasoning process. Particularly, for a specific state, the inspectable bottleneck structure of SCoBots allows to pinpoint not just the object properties, but importantly the relational concepts being used for the final action decision. This is exemplified in Fig. 4 on Skiing, Pong and Kangaroo (cf. App. A.4 for explanations of SCoBots on the remaining games). Here we highlight the decision-making path (from SCoBots’ decision tree based policies), at specific states of each game. For example, for the specific game state extracted from Skiing, the SCoBot agent selects RIGHT as the best action, because the signed distance from its character to the left flag is larger than a specific value (+15+15 pixels). Given the nature of the game this inherent explanation suggests that the agent is indeed basing its selection process on relevant relational concepts.
@@
A more striking display of the benefits of the inherent explanations of SCoBot agents is depicted by the Pong agent in Fig 4. Specifically, the provided explanation suggests that the agent is basing its decision on the vertical positions of the enemy and of its own paddle (distance between the two paddles on the $y$ axis). In fact, this suggests that the agent is largely ignoring the ball’s position. Interestingly, upon closer inspection of the enemy movements, we observed that, previously unknown, the enemy is programmed to follow the ball’s vertical position (with a small delay). Thus, the vertical positions of these two objects are highly correlated (Pearson’s and Spearman’s correlations coefficient above $99\%$, cf. App. A.7). Moreover, the ball’s rendering can lead to visual flickering, explaining why SCoBots base their decision on the potentially more reliable feature of the enemy paddle’s vertical position.

### v4

@@ 1 Introduction
Deep Reinforcement learning (RL) agents are prone to suffer from a variety of issues that hinder them from learning optimal or generalizable policies. Prominent examples are reward sparsity (Andrychowicz et al., 2017) and difficult credit assignment (Raposo et al., 2021; Wu et al., 2024). A more pressing issue is the goal misalignment problem. It occurs when an agent optimizes a different side-goal, aligned with the original target goal during training (Koch et al., 2021), but not at test time. Such misalignments can be difficult to identify (di Langosco et al., 2022). For instance, in this work we discover that such a misalignment can occur in the oldest and most iconic video game, Pong (cf. Fig. 1). In Pong, the agent’s target goal is to catch a ball with its own paddle and to return it passed the enemy’s one. The enemy is programmed to constantly follow the ball, thus, agents can learn to focus on the position of the enemy paddle for placing their own, rather than the position of the ball itself. If left unchecked, such shortcut learning (Geirhos et al., 2020) may lead to a lack of model generalization and unintuitive failures e.g. at deployment time (Zhang et al., 2021).
@@
Figure 1: Successive Concept Bottlenecks Agents (SCoBots) allow for easy inspection and revision. Top: Deep RL agents trained on Pong produce high playing scores with importance map explanations that suggest sensible underlying reasons for taking an action ( Refer to caption ). However, when the enemy is hidden, the deep RL agent fails to even catch the ball without clear reasons ( Refer to caption ). Bottom: SCoBots, on the other hand, allow for multi-level inspection of the reasoning behind the action selection, e.g., at a relational concept, but also action level. Moreover, they allow users to easily intervene on them ( Refer to caption ) to prevent the agents from focusing on potentially misleading concepts. In this way, SCoBots can mitigate RL specific caveats like goal misalignment.
@@
Recently eXplainable AI (XAI) methods have emerged to detect such shortcut behavior, by identifying the reasons behind a model’s decisions (Schramowski et al., 2020; Roy et al., 2022; Saeed and Omlin, 2023). However, many XAI methods’ explanations do not faithfully present the model’s underlying decision process (Chan et al., 2022). For example, importance-map explanations indicate the importance of an input element without indicating why this element is important (Kambhampati et al., 2021; Stammer et al., 2021; Teso et al., 2023). A recent branch of research therefore focuses on models that provide inherent concept-based explanations. Prominent examples are concept bottlenecks models (CBMs), which provide predictive performances on par with standard deep learning approaches for supervised image classification (Koh et al., 2020). More importantly, CBMs allow to identify and revise incorrect model behavior on a concept level (Stammer et al., 2021).
@@
Contrary to exising CBMs’ fields of applications, RL requires relational reasoning (Kaiser et al., 2019). In this work, we therefore introduce Successive Concept Bottleneck Agents (SCoBots, cf. Fig. 1, bottom) bringing the concept bottleneck approach to RL. SCoBots integrate successive concept bottlenecks into their decision processes, where each bottleneck layer provides concept representations that integrate the concept representations of the previous bottleneck layer. Specifically, provided a set of predefined concept functions, SCoBots automatically extract relational concept representations based on objects and their properties of the initial bottleneck layers. Finally, the set of relational and object concept representations are used for optimal action selection. SCoBots thus represent inherently explainable RL agents that, in comparison to deep RL agents, allow for inspecting and revising their learned decision policies at multiple levels of their reasoning processes: from single object properties, through relational concepts to the action selection.
@@
Our evaluations on the iconic Atari Learning Environments (ALE, Mnih et al. (2013)) provides experimental evidence that SCoBots perform on par with deep RL agents. More importantly, we showcase SCoBots’ ability to provide valuable explanations and the potential of mitigating a variety of RL specific issues, from reward sparsity to misalignment problems, via simple guidance from domain experts. We identify previously unknown shortcut behavior of deep agents, even on the simple Pong game. By utilizing the interaction capabilities of SCoBots, this behavior is easily corrected. Ultimately, our work illustrates the severity of goal misalignment issues in RL and the importance of being able to mitigate these and other RL specific issues, via relational concept based models.In summary, our contributions are:
(i) We introduce Successive Concept Bottleneck agents (SCoBots).
@@
(ii) We show that SCoBots allow to inspect their internal decision processes.
@@
(iii) We show that their inherent inspectable nature can helps identifying unknown misalignments.
@@
(iv) We show that they allow for human interactions for mitigating various RL specific issues.
@@
We proceed as follows. We introduce SCoBots and discuss their specific properties. We continue with experimental evaluations and analysis. Before concluding, we touch upon related work.
@@
Figure 2: An overview of Successive Concept Bottlenecks Agents (SCoBots). SCoBots decompose the policy into consecutive interpretable concept bottlenecks (ICB). Objects and their properties are first extracted from the raw input, human-understandable functions are then employed to derive relational concepts, used to select an action. The understandable concepts enable interactivity. Each bottleneck allows expert users to, e.g., prune or utilize concepts to define additional reward signals.
@@ 2 Successive Concept Bottleneck Agents
In this work, we represent RL problems through the framework of a Markov Decision Process, $\mathcal{M}=<\mathcal{S},\mathcal{A},P_{s,a},R_{s,a},\gamma>$, with $\mathcal{S}$ as the state space, $\mathcal{A}$ the set of available actions, $P(s,a)$ the transition probability, and $R(s,a)$ the immediate reward function, obtained from the environment, and $\gamma$ the didscount factor. Classic deep RL policies, $\pi_{\theta}(s)=P(A=a|S=s)$ parameterized by $\theta$, are usually black-box models that process raw input states, e.g. a set of frames, to provide a distribution over the action space (Mnih et al., 2015; van Hasselt et al., 2016).
@@
The main body of research on CBMs focuses on image classification tasks, where the extracted concepts represent attributes of objects. In contrast, RL agents must learn not just from static data, but also through interaction with dynamically evolving environments. RL tasks thus often require relational reasoning, as they involve understanding the relationships between instances that evolve through time and interact with another. This is crucial for learning effective policies in complex, dynamic environments. Note that, in the following descriptions, we use specific fonts to distinguish objects’ properties from their relations .
@@ 2.1 Building inspectable ScoBots
An underlying assumption of SCoBots is that their processing steps should be inspectable and understandable by a human user. This stands in contrast to e.g. unsupervised CBM approaches Jabri et al. (2019); Zhou et al. (2019); Srinivas et al. (2020), where there is no guarantee for learning human-aligned concepts. SCoBots rather take a different approach by dividing the concept learning and RL-specific action selection into several inspectable steps that are grounded in human-understandable concept representations. These steps are described hereafter and depicted in Fig. 2.
@@
One of the major differences to previous work on CBMs is that SCoBots further extract and utilize relational concepts that are based on the previously extracted objects and their properties. Formally, SCoBots utilize a consecutive bottleneck model, $\mu_{\mathcal{F}}(\cdot)$, called the relation extractor. This model, $\mu$, is parameterized by a predetermined set of transparent relational functions, $\mathcal{F}$, used to extract a set of $d_{t}$ relational concepts. These relations are based on each individual object (and its properties) for unary relations or on combinations of objects for n-ary relations, and are denoted as $\mu_{\mathcal{F}}(\Omega_{t})=\Gamma_{t}=\{g_{t}^{k}\}_{k=1}^{d_{t}}$. Without strong prior knowledge, $\mathcal{F}$ can initially correspond to universal object relations such as distance and speed . However, this set can easily be updated on the fly by a human user with e.g. additional, novel relational functions. Note that $\mathcal{F}$ can include the identity, to let the relation extractor pass through initial object concepts from $\Omega_{t}$ to the relational concepts $\Gamma_{t}$.
@@
Finally, SCoBots employ an action selector, $\rho_{\theta_{2}}$, parameterized by $\theta_{2}$}, on the relational concepts to select the best action, $a_{t}$, given the initial state, $s_{t}$ (i.e. the set of frames). Up to now, all extracted concepts in SCoBots, both $\Omega_{t}$ and $\Gamma_{t}$, represent human-understandable concept representations. To guarantee understandability also within the action selection step of SCoBots, we need to embed an interpretable action selector. While neural networks, a standard choice in RL literature, are performative and easy to train using differentiable optimization methods, they lack this required interpretability feature. However, Çağlar Aytekin (2022) have recently shown the equivalence between ReLU-based neural networks and decision trees, which in contrast represent inherently interpretable models. To trade-off these issues of flexibility and performance vs interpretability, SCoBots thus break down the action selection process by initially training a small ReLU-based neural network action selector, $\widetilde{\rho}_{\theta^{\prime}_{2}}$, via gradient-based RL optimization. After this, $\widetilde{\rho}_{\theta^{\prime}_{2}}$ is finally distilled into a decision tree $\rho_{\theta_{2}}$. Lastly, note that one can add a residual link from the initial object concepts, $\Omega_{t}$, to the relational concepts, $\Gamma_{t}$ such that the action selector can, if necessary, also make decisions based on basic object properties, e.g., the height of an object.
@@
Overall, our approach preserves the MDP formulation used by classic deep approaches, but decomposes the classic deep policy $s_{t}\xrightarrow{\pi_{\theta}}a_{t}$ into a successive concept bottleneck one $s_{t}\xrightarrow{\omega_{\theta_{1}}}\Omega_{t}\xrightarrow{\mu_{\mathcal{F}}}\Gamma_{t}\xrightarrow{\rho_{\theta_{2}}}a_{t}$, where $\theta=(\theta_{1},\theta_{2},\mathcal{F}$) constitutes the set of policy parameters. For simplicity, we will discard the parameter notations in the rest of the manuscript. Further explanations of the input and output space of each module, as well as the properties and relational functions used in this work are provided in the appendix (cf. App.A.5 and App. A.6).
@@
Instead of jointly learning object detection, concept extraction, and policy search, SCoBots enable independent optimization of each policy component. Separating the training procedure of different components reduces the complexity of the overall optimization problem (Koh et al., 2020).
@@ 2.2 Guiding SCoBots
Concept pruning. The type and amount of concepts that are required for a successful policy may vary across tasks. For instance, the objects colors are irrelevant in Pong but required to distinguish vulnerable ghosts from dangerous ones in MsPacman. However, overloading the action selector with irrelevant concepts, whether these are object properties ($\Omega_{t}$) or relational concepts ($\Gamma_{t}$), can lead to difficult optimization (e.g. the agent focusing on noise in unimportant states) as well as difficult inspection of the decision tree ($\rho$). Moreover, for a single RL task, the need for specific concepts might even change during training, in e.g. progressive environments (where agents need to master early stages before being provided with additional, more complex tasks (Delfosse et al., 2024c)).
@@
For example, one can impose penalties, based on the distance between the agent and objects (cf. “expert reward” arrow in Fig. 2), with the intention of incentivizing the agent to maintain close proximity with the ball. As shown in our experimental evaluation, we can use this concept pruning and concept based reward shaping to easily address many RL specific caveats such as reward sparsity, ill-defined objectives, difficult credit assignment, and misalignment (cf. App. A.8 for details on each problem).
@@ 3 Experimental Evaluations
In our evaluations, we investigate several properties and potential benefits of the transparent SCoBot agents. We specifically aim to answer the following research questions:
@@
(Q2) Does the inspectable nature of SCoBots allow to detect issues in their decision processes?
@@
(Q3) Can concept-based guidance help mitigate common RL caveats, such as policy misalignments?
(Q4) Can SCoBots learn with imperfect object extraction methods?
(Q5) How crucial is the the relation extractor for SCoBots performances?
@@
Experimental setup: We evaluate SCoBots on $9$ Atari games (cf. Fig. 3 from the Atari Learning Environments (Bellemare et al., 2012) (by far the most used RL framework (cf. App. A.1), as well as the HackAtari modified (Delfosse et al., 2024a) Pong environments, where the enemy is not visible yet active (NoEnemy), and where the enemy stops moving after returning the ball (LazyEnemy). We provide human normalized scores (following eq. 3) that are averaged over $3$ seeds for each agent configuration. We compare our SCoBot agents to deep agents with the classic convolutional network introduced by Mnih et al. (2015), that process a stack of $4$ black and white frames and denote these as deep agents in the following. Note, that we evaluate all agents on the latest v5 version of the environments [^2] This leads to deep agents overall slightly worse than in Schulman et al. (2017), however we obtain similar results when evaluating on the old v4 versions (cf. Tab. 1). to prevent overfitting. All agents are trained for $20\text{M}$ frames under the Proximal Policy Optimization algorithm (PPO, (Schulman et al., 2017)), specifically the stable-baseline3 implementation (Raffin et al., 2021) and its default hyperparameters (cf. Tab. 2 in App. A.5).
@@
We focus our SCoBot evaluations on the more interesting aspects of the agent’s reasoning process underlying the action selection, rather than the basic object identification step (which has been thoroughly investigated in previous works e.g. (Koh et al., 2020; Stammer et al., 2021; Wu et al., 2024)). We thus assume access to a pretrained object extractor and provide our agents with object-centric descriptions of these states ($\Omega_{t}$) based on the information from OCAtari (Delfosse et al., 2024b). Specifically, in our evaluations, the set of object properties consists of object class (e.g. enemy paddle, ball etc.), $(x^{t},y^{t})$ coordinates, coordinates at the previous position $(x^{t-1},y^{t-1})$, height and width, the most common RGB values (i.e. the most representative color of an object), and object orientation. The specific set of functions, $\mathcal{F}$, that are used to extract object relations is composed of: euclidean distance, directed distances (on $x$ and $y$ axes), speed, velocity, plan intersection, the center (of the shortest line between two objects), and color name (cf. App. A.5.1 for implementation details).
@@
To distill SCoBots’ learned policies, we use the decision-tree extraction algorithm VIPER (Bastani et al., 2018). For the human-guided SCoBot evaluations (denoted as (guided) SCoBots in our evaluations), we illustrate human guidance and prune out the concepts that we consider unimportant to master the game (cf. Tab. 4). Furthermore, to mitigate RL specific problems, we also provide SCoBot agents with simulated human-expert feedback signals, the details of which we describe at the relevant sections below. More details about the setup and the hyperparameters are in App. A.5. In our evaluation, SCoBots’ training is slightly faster than deep agents’ one (cf. Appendix A.9).
Figure 3: Object-centric agents can master different Atari environments and interactive SCoBots allow for corrections. Human-normalized scores of different agents trained using PPO on 99 ALE environments, including deep agents (i.e. using CNNs), guided decision tree policy (SCoBots), their neural object-centric baseline (NN-), and these baselines without guidance (NG). SCoBots obtain similar or better scores than the deep agents, showing that object-centric agents can also solve RL tasks while making use of human-understandable concepts (left). Guiding SCoBots allow to correct misalignment in Pong (center) and to obtain the originally intended agents, depicted by a level completion score of 100% on the intended goal’s evaluation in Kangaroo (right).
@@
Figure 4: Interpretable SCoBots allow to follow their decision process, thanks to their interpretable concepts. The states and associated decision processes of SCoBots (extracted from the decision trees) on Skiing (left), and from unguided SCoBots on Pong (middle) and Kangaroo (right). For example, in this Skiing state, our SCoBot selects RIGHT, as the signed distance between Player and the (left) Flag1 (on the $x$ axis) is bigger than $15$. This agent selects the correct action for the right reasons.
@@
The main target of developing ScoBots is to obtain competitive, yet transparent agents that allow for human users to identify their underlying reasoning process. Particularly, for a specific state, the inspectable bottleneck structure of SCoBots allows to pinpoint not just the object properties, but importantly the relational concepts being used for the final action decision. This is exemplified in Fig. 4 on Skiing, Pong and Kangaroo (cf. App. A.4 for explanations of SCoBots on the remaining games). Here we highlight the decision-making path (from SCoBots’ decision tree based policies), at specific states of each game. For example, for the game state extracted from Skiing, the SCoBot agent selects RIGHT as the best action, because the signed distance from its character to the left flag is larger than a specific value (+15+15 pixels). Given the nature of the game this inherent explanation suggests that the agent is indeed basing its selection process on relevant relational concepts.
@@
A more striking display of the benefits of the inherent explanations of SCoBot agents is depicted by the Pong agent in Fig 4. The provided explanation suggests that the agent is basing its decision on the vertical positions of the enemy and of its own paddle (distance between the two paddles on the $y$ axis). In fact, this suggests that the agent is largely ignoring the ball’s position. Interestingly, upon closer inspection of the enemy movements, we observed that, previously unknown, the enemy is programmed to follow the ball’s vertical position (with a small delay). Thus, the vertical positions of these two objects are highly correlated (Pearson’s and Spearman’s correlations coefficient above $99\%$, cf. App. A.7). Moreover, the ball’s rendering contains flickering, explaining why SCoBots base their decision on this potentially more reliable feature, the enemy paddle’s vertical position.